Rain appears to be fallin at at angle of 37 degree with vertical to the driver of a car moving with a velocity 7m/sec. Whwn he increases the velocity of the car to 25 m/sec, the rain again appears to fall at an angle 37 degree wjth vertical. What is the actual velocity of rain and its direction with vertical?

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Pooja,

Solve for b and c first:

Horizontal Vertical
s = ?
u = 75.2cos34.5= 61.9743 m/s
v = "
a = 0
t = ?
s = 0
u = 75.2sin34.5= 42.5937 m/s
v = -42.5937 m/s (hitting level ground)
a = -9.8 m/s/s
t = ?
  Use:
v = u + at
-42.5937 = 42.5937 + -9.80t
t = 8.6926 s = 8.69 s (b)
Use:
vav = Ds/Dt
61.9743 m/s = Ds/8.6926 s
s = 538.7 m = 539 m (c)
 

now solve a:

Horizontal Vertical
s = ?
u = 75.2cos34.5= 61.9743 m/s
v = "
a = 0
t = ?
s = ????
u = 75.2sin34.5= 42.5937 m/s
v = 0 (at the top)
a = -9.8 m/s/s
t = ?
  Use:
v2 = u2 + 2as
02 = (42.5937)2 + 2(-9.8 m/s/s)s
s = 92.56 m = 92.6 m (a)

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert